OpenStudy (anonymous):

Ed is doing a survey of popular colors for cars for a school report. He decides to take a sample by sitting near a busy intersection and noting the colors of the cars as they go by him. Ed counts 23 black cars, 9 red cars, 17 blue cars, 25 white cars, and 21 silver cars during 20 minutes of watching. How many more white and blue cars are there than silver and red cars?

OpenStudy (anonymous):

Once again, add both blue and white and subtract with silver and red. Equation: 17+25=42 21+9=30 42-30=8 8 is your answer.

OpenStudy (anonymous):

Thanks, but I got 9